# Expert-Token Resonance MoE: Bidirectional Routing with Efficiency Affinity-Driven Active Selection

### v3

@@ Introduction
The proportional time consumption at the operator level is depicted in Figure 6. Among the components, AI CORE efficiently executes matrix multiplications and convolutions in AI algorithms; AI VECTOR CORE accelerates vector operations through parallel processing; MIX AIC integrates different types of operators and optimizes for multiple tasks; AI CPU is optimized in hardware and instruction sets to better support AI algorithms. Our approach selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE. This leads to an overall 2.8$\times$ reduction in the cumulative time consumption of the MatMul operator and a 2.6$\times$ decrease in Cube computing load. However, the proportions of TopK and IndexPutV2, related to rearrangement, show a slight increase.
Figure 6: The distribution of time consumption for operators.
@@
We select a single iteration during the stable training period and describe the per-device memory usage (Allocated) using the first 100,000 samples from its memory monitoring, as shown in Figure 5. Overall, our approach achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE. As cluster size increases, the proportion of computational overhead decreases, and the gap in memory usage narrows. Additionally, instantaneous memory peaks gradually disappear, and the fluctuation amplitude of short-term memory also diminishes.
The Performance of Downstream Tasks
To enhance the model’s conversational capabilities and adaptability to downstream task, we fine-tuned the pre-trained models. As shown in Figure 7, with sufficient supervised fine-tuning (SFT), our approach achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE. The Rewriting and Summary capabilities show the highest improvement, with a 28.2% increase compared to the Baseline and a 6.7% increase compared to LocMoE. In the 13 tests of Domain Competency Exam, our approach demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE. The IP Training in the digital communications domain shows the most significant improvement, with a 27.3% increase compared to the Baseline and a 3.0% increase compared to LocMoE. Among the 18 sub-capabilities of General Ability, our approach exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE. The capability of Planning demonstrates the highest improvement, with a 26.8% increase compared to the Baseline and a 2.92% increase compared to LocMoE.
Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly. Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 13.6% compared to the Baseline and 2.8% compared to LocMoE in the GPQA (Rein et al. 2023) evaluation, despite the limited data available for training. During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies. TeleQnA (Maatouk et al. 2023), the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain. Consequently, our approach comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
Table 1: Performance promotion obtained by our approach on different datasets.
GDAD
GDAD-1 GDAD-2 GDAD-3 Avg GPQA TeleQnA
Baseline 47.8 43.0 65.4 52.8 29.5 62.1
LocMoE 55.5 47.6 71.1 59.0 32.6 67.6
LocMoE+ 57.4 49.9 74.5 61.5 33.5 68.8
@@ Conclusion
In this paper, we propose a novel expert routing framework that enhances MoE efficiency through three key innovations: an efficient routing mechanism with lightweight computation, a bidirectional expert-token resonance selection mechanism, which combined ECR and TCR, and a dynamic capacity bounds module. The framework integrates orthogonal feature extraction and optimized expert localization loss, effectively addressing expert homogeneity while improving routing performance. Our local expert strategy demonstrates advantages in both load balancing and communication efficiency. Experimental results validate the effectiveness of the proposed framework across multiple benchmarks. Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively. To evaluate model performance, all models are evaluated with the open-source datasets GPQA and TeleQnA, and closed domain benchmark GDAD. In downstream tasks, our approach outperforms the Baseline by 14.1%, 13.6%, and 9.7% on GDAD, GPQA, and TeleQnA, respectively. Future work may explore methods to compress communication data to further reduce communication overhead.
@@
$=\sum_{i=1}^{n}\mathcal{P}(\bm{o}_{i}\text{ sent to expert }i|\bm{o}_{i}\text{ is in }\bm{x})\cdot\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
@@
$=\sum_{i=1}^{n}\mathcal{P}(\text{expert }i\text{ choose }\bm{o}_{i}|\bm{o}_{i}\text{ is in }\bm{x})\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
@@
Datasets for Training and Fine-Tuning
@@
The experiments are conducted on a cluster composed of Ascend 910B3 NPUs, divided into three groups: 32 NPUs (hereinafter referred to as 32N, and so on), 64N, and 256N. The 910B3 series NPU contains 20 AI cores with a main frequency of 1.8GHz and a theoretical computing power of 313T under fp16 precision. The physical High Bandwidth Memory (HBM) of the 910B3 NPU is 64G, with an HBM frequency of 1.6GHz and an HBM bandwidth of 1.6T. Every 8 NPUs are mounted on the same Atlas 800T A2 server, which internally adopts a fullmesh networking scheme, meaning that any two NPUs are interconnected. The version of the Ascend Hardware Development Kit (HDK) is 23.0.2.1, and the version of the Compute Architecture for Neural Networks (CANN) suite is 7.0.0, which is the commercial release version for Q4 2023. The models in this paper use ModelLink, an LLM training framework based on the Ascend architecture, and run in the torch_npu 5.0.0 environment.
Model Configuration
This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and our approach. Mixtral has 46.7B parameters and uses the Group Query Attention (GQA) mechanism to compute attention. It contains 32 sparse expert blocks, with the MoE Feedforward layer comprising 8 experts, and each token selects the top 2 experts for processing. In the application scenarios of this paper, the corpora are generally long texts; thus, the sequence length is adjusted to 32768. For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2. The batch size is set to 128.
@@
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs C-Eval (Huang et al. 2024b) and TeleQnA (Maatouk et al. 2023) to evaluate the model’s Chinese language capability.

### v4

@@ Introduction
Expert Homogenization and Load Distribution Analysis
The Calinski-Harabasz (CH) index (Lima and Cruz 2020) measurements reveal that bidirectional affinity selection significantly enhances token clustering quality in MoE architectures:
[…]
While baseline and single-mechanism approaches achieve comparable improvements, the integrated LocMoE+ method demonstrates superior performance, as shown in Figure 9. Combining token-to-expert and expert-to-token selection mechanisms creates synergistic effects that accelerate natural clustering tendencies during training. Our bidirectional approach establishes a positive feedback loop between token routing and expert affinity, fundamentally enhancing expert utilization efficiency through improved feature organization and specialization.
Figure 9: The Calinski-Harabasz index across training steps.
Figure 10 reveals fundamental differences in how various loss functions affect token distribution across experts. The baseline approach suffers from severe load imbalance, with certain experts becoming overloaded while others remain idle. The auxiliary loss method provides marginal improvements through regularization, yet distribution remains significantly skewed. The locality loss demonstrates transformative effects by incorporating architectural topology into the optimization objective, achieving balanced token allocation across all experts through KL divergence constraints that simultaneously minimize inter-node communication and prevent expert collapse.
Figure 10: The distribution of tokens assigned to experts with different loss function.
Figure 11 presents the cumulative distribution function (CDF) and empirical cumulative distribution function (ECDF) analysis across these routing methods throughout the training progression. The locality loss approach presents distinctly optimal characteristics across both CDF and ECDF measurements, maintaining consistently high performance levels throughout the training process with remarkable stability during later training phases. The sustained performance across different probability and load thresholds indicates that incorporating expert-token affinity into the routing objective creates robust optimization dynamics that preserve both routing quality and load distribution efficiency. These findings underscore the effectiveness of ETR in addressing the inherent challenges of expert-token assignment optimization, providing a principled foundation for scalable sparse model.
Figure 11: The CDF and ECDF of different schemes.
The Performance of Downstream Tasks
The GDAD benchmark, comprising three distinct evaluation tasks, serves as the primary assessment framework for domain task capabilities, while TeleQnA provides specialized validation for telecommunications knowledge applications. Our integrated LocMoE+ approach consistently outperforms both baseline and single-mechanism variants across all evaluation metrics (see Table 1), achieving notable gains through bidirectional selection mechanisms. The consistent improvements observed across both general domain tasks and specialized telecommunications applications confirm that the synergistic combination of bidirectional selection mechanisms creates substantial performance advantages while validating the robustness and generalizability of our proposed architecture modifications.
Table 1: Domain performance promotion obtained by our approach on different datasets.
GDAD
GDAD-1 GDAD-2 GDAD-3 Avg TeleQnA
Baseline 47.8 43.0 65.4 52.8 62.1
LocMoE (TCR) 55.5 47.6 71.1 59.0 67.6
LocMoE (ECR) 45.8 45.6 62.8 56.3 61.8
LocMoE+ 57.4 49.9 74.5 61.5 68.8
Table 2 presents general performance evaluation results across three widely-recognized benchmarks—MMLU (Hendrycks et al. 2021) for comprehensive knowledge assessment, GPQA (Rein et al. 2023) for advanced reasoning capabilities, and HumanEval (Chen et al. 2021) for code generation proficiency—revealing distinct performance characteristics of our proposed methods. The results demonstrate that our bidirectional LocMoE+ approach achieves superior performance in reasoning and coding tasks while maintaining competitive general knowledge capabilities, with individual constraint routing mechanisms exhibiting complementary strengths across different evaluation dimensions. While the baseline maintains slight advantage in MMLU, the integrated LocMoE+ approach demonstrates that bidirectional selection mechanisms create meaningful improvements in task-specific capabilities without substantial degradation in general knowledge retention, suggesting that our architectural modifications enhance model specialization for complex reasoning and generation tasks while preserving foundational knowledge capabilities.
Table 2: General performance comparison of different MoE methods
Method MMLU GPQA HumanEval
Baseline 71.8 29.2 40.2
LocMoE (TCR) 68.4 30.3 52.8
LocMoE (ECR) 45.8 32.5 57.6
LocMoE+ 70.4 33.5 67.8
To enhance conversational capabilities and downstream task adaptability, we conducted supervised fine-tuning on the pre-trained models. As shown in Figure 12, our approach demonstrates substantial improvements across multiple evaluation dimensions within the General and Domain-specific Assessment Dataset (GDAD). The method achieves an average improvement of approximately 20.1% across 16 sub-capabilities of Domain Task Capability compared to the baseline, with particularly notable gains in rewriting and summary capabilities. In the Domain Competency Exam assessments, our approach shows an average improvement of 16% relative to the baseline, with IP Training in digital communications demonstrating the most significant advancement. Among the 18 sub-capabilities of General Ability, the method exhibits an improvement of about 13.9% relative to the baseline, with planning capabilities showing the highest enhancement at 26.8%.
@@ Conclusion
In this paper, we propose ETR, a fundamentally new approach to MoE routing that solves the longstanding trade-off between computational efficiency and model performance through theoretically-grounded bidirectional selection mechanisms. By dynamically coordinating token-choice and expert-choice routing based on training progress, ETR achieves simultaneous improvements in both training efficiency and downstream task quality. The substantial performance gains demonstrated across diverse benchmarks, combined with significant reductions in computational overhead, establish ETR as a critical advancement for practical deployment of large-scale sparse models. Our theoretical contributions provide new insights into expert-token dynamics that extend beyond incremental optimizations, opening pathways for next-generation MoE architectures.
@@
Datasets for Pre-training and Fine-Tuning
@@
The pre-training data comprises 300B tokens in total, with 150B tokens from the ICT domain and 150B tokens sampled from general data. The sampling ratios are shown in Table 3. For SFT data, we employed a two-stage training: the first stage primarily enhances the model’s logical reasoning capabilities such as multi-task capability, mathematics, puzzle-solving, complex logic, etc. The total scale of samples is approximately two million, while the second stage focuses on improving instruction-following abilities, tool function call, sentiment, security, etc. The total scale of samples in stage2 is about three million.
Table 3: Data sources and sampling ratios of general pre-training data.
Primary Category Secondary Category Tertiary Source Sampling ratio
General English Webpages Reasoning_steplist 25%
Model_rewrite 100%
Books & Papers book3 25%
bookcorpus 100%
all_libgen_books 20%
all_libgen_scihub 10%
RedPajama_arxiv 25%
arxiv_latex2Markdown_cleaned 25%
WebText wiki 100%
stackexchange_cleaned 20%
cosmopedia_v2 15%
General Chinese Webpages aigc_dataset 15%
Book all_book_deduped 10%
zh_book_CommonData 10%
zh_general_STEM 80%
all_zhiwang 20%
WebText baike_MBAzhiku_sougou_ye_zhiarge 50%
baike_sougou_baidu_kuaidong 50%
wiki 10%
zhihu_caigou_merged_cleaned 10%
High-density Knowledge Q&A quiz_data 100%
Collection density_knowledge 100%
collection_updated 100%
english_question_and_answer 100%
annealing 100%
Code Code code_python_edu_high_quality 30%
Forum CSDN 20%
Ultra_textbooks 100%
@@
The experiments are conducted on a cluster composed of Ascend 910B3 NPUs, divided into three groups: 32 NPUs (hereinafter referred to as 32N, and so on), 64N, and 256N. The 910B3 series NPU contains 20 AI cores with a main frequency of 1.8GHz and a theoretical computing power of 313T under fp16 precision. The physical High Bandwidth Memory (HBM) of the 910B3 NPU is 64G, with an HBM frequency of 1.6GHz and an HBM bandwidth of 1.6T. Every 8 NPUs are mounted on the same Atlas 800T A2 server, which internally adopts a fullmesh networking scheme, meaning that any two NPUs are interconnected.
@@
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs GPQA (Rein et al. 2023) and TeleQnA (Maatouk et al. 2023) to evaluate the model’s Chinese language capability.